What is the pH of solutions having the following H3O+ concentrations? Identify each as acidic, basic, or neutral.

  1. (a) [H3O+] = 1.0 × 10−3M

  2. (b) [H3O+] = 1.0 × 10−13M

  3. (c) [H3O+] = 3.4 × 10−10M
Answer #1

Acidic solutions have pH less than 7

Neutral solutions have pH = 7

Basic solutions have pH greater than 7

let us calculate pH of each solution

pH = -log [H3O+]

a) [H3O+] = 1.0 × 10−3M

pH = -log(1x10^-3) = 3

Since pH <7 , solution is acidic
**********************
(b) [H3O+] = 1.0 × 10−13M
pH = -log [H3O+]
pH = -log(1x10^-13) = 13

Since pH >7 , solution is basic
*****************************

(c) [H3O+] = 3.4 × 10−10M

pH = -log [H3O+]
pH = -log(3.4x10^-10) = 9.47

Since pH >7 , solution is basic
*******************************

kindly upvote .Thank you